OpenStudy (anonymous):

find four consecutive integers if the sum of the third and fourth is 63

OpenStudy (lgbasallote):

x+3 + x+4 = 63 x = 28 x+1 = 29 x+2 = 30 x+3 = 31 x+4 = 32